OpenStudy (volpina):

What, if anything, does the Mean Value Theorem guarantee for the given function on the given interval? f(x)=x^2-2x+5 on [1,4]

OpenStudy (jiteshmeghwal9):

Find f(1)=f(4) if they r equal then there is atleast one number c in the interval [1,4] for which f(c)=0

OpenStudy (mathmale):

I'd suggest you review your textbook's presentation of the M. V. T. This Theorem involves an equation, but also two conditions that must be satisfied for the M. V. T.'s conclusion to be valid. Focusing on the equation: It states that the slope of the line connecting the end points of your graph on the given interval equals the derivative of your function evaluated at some number c. Your job is to find the vaue of c. \[\frac{ f(b)-f(a) }{ b-a }=f '(c)\]

OpenStudy (mathmale):

Condition 1 for the MVT: your f(x) must be continuous on the given interval. Condition 2 : your f(x) must be differerentiable on that interval. Are these conditions met?

OpenStudy (3mar):

The Mean Value Theorem states that if f(x) is \(\color{red}{defined} \) and \(\color{red}{continuous} \) on the interval [a,b] and \(\color{red}{differentiable} \) on (a,b), then there is at least one number c in the interval (a,b) (that is a < c < b) such that: \[\LARGE f '(\color{red}c)=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (mathmale):

volpina: Please involve yourself in this discussion.

OpenStudy (volpina):

So do i have to substitute into the Mean Value Theorm's equation and see if it meets the conditions?

OpenStudy (mathmale):

There's more to it than that. FIRST, you determine whether or not the 2 conditions are satisfied. Then, you calculate the quantities on each side of the given equation and set them equal to one another. As before, your job is to determine whether nor not there is some value c between x=a and x=b at which this equation is true. "What, if anything, does the Mean Value Theorem guarantee for the given function on the given interval?" Reread the MVT and focus on the CONCLUSION.

OpenStudy (mathmale):

\[\LARGE f '(\color{red}c)=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (mathmale):

You may as well go ahead and calculate all of the components in the above equation. Given a and b values, evaluate f(a) and f(b). Then calculate \[\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (mathmale):

May I see your work and your result, please?

OpenStudy (volpina):

f(4)=(4)^2-2(4)+5=13 f(1)=(1)^2-2(1)+5=4 13-4/4-1 9/3 3

OpenStudy (mathmale):

OK. Now calculate \[\frac{ f(b)-f(a) }{ b-a }\] I see y ou've already given that a stab, but you haven't labeled your results.

OpenStudy (mathmale):

Is this what you meant?\[\frac{ f(b)-f(a) }{ b-a }=\frac{ 13-4 }{ 4-1 }?\]

OpenStudy (mathmale):

If so, calculate (and label) it.

OpenStudy (volpina):

\[\frac{ f(b)-f(a) }{ b-a }=\frac{ 13-4 }{ 4-1 }=3?\]

OpenStudy (mathmale):

Now please find the derivative of the given function. f '(x)= ??

OpenStudy (volpina):

from the original function, it would 2x-2?

OpenStudy (mathmale):

Yes. f '(x) = 2x-2. Now replace both "x" with "c." Write out your result.

OpenStudy (mathmale):

volpina?

OpenStudy (volpina):

is it 4?

OpenStudy (mathmale):

Your derivative was f '(x) = 2x - 2. I'm asking you to throw out both "x" and replace them with 'c.' please do that now. f '(c) = ?

OpenStudy (mathmale):

We need to follow the road map of this Theorem closely.

OpenStudy (mathmale):

We have found [f(b)-f(a)]/[b-a] and now have to find f '(c).

OpenStudy (mathmale):

You do not yet have a value for c. Instead, you have an equation for c.

OpenStudy (mathmale):

volpina, where have you been?

OpenStudy (volpina):

f '(c) = 2c - 2.